# Generalized Out-of-Distribution Detection: A Survey

### v1

@@ Abstract
Out-of-distribution (OOD) detection is critical to ensuring the reliability and safety of machine learning systems. For instance, in autonomous driving, we would like the driving system to issue an alert and hand over the control to humans when it detects unusual scenes or objects that it has never seen before and cannot make a safe decision. This problem first emerged in 2017 and since then has received increasing attention from the research community, leading to a plethora of methods developed, ranging from classification-based to density-based to distance-based ones. Meanwhile, several other problems are closely related to OOD detection in terms of motivation and methodology. These include anomaly detection (AD), novelty detection (ND), open set recognition (OSR), and outlier detection (OD). Despite having different definitions and problem settings, these problems often confuse readers and practitioners, and as a result, some existing studies misuse terms. In this survey, we first present a generic framework called generalized OOD detection, which encompasses the five aforementioned problems, i.e., AD, ND, OSR, OOD detection, and OD. Under our framework, these five problems can be seen as special cases or sub-tasks, and are easier to distinguish. Then, we conduct a thorough review of each of the five areas by summarizing their recent technical developments. We conclude this survey with open challenges and potential research directions.
@@
Anomaly Detection, Novelty Detection, Open Set Recognition, Out-of-Distribution Detection, Outlier Detection
@@ 1 Introduction
Atrustworthy visual recognition system should not only produce accurate predictions on known context, but also detect unknown examples and reject them (or hand them over to human users for safe handling) [1, 2, 3, 4, 5]. For instance, a well-trained food classifier should be able to detect non-food images such as selfies uploaded by users, and reject such input instead of blindly classifying them into existing food categories. In safety-critical applications such as autonomous driving, the driving system must issue a warning and hand over the control to drivers when it detects unusual scenes or objects it has never seen during training.
Figure 1: Taxonomy of generalized OOD detection framework. Four bases are used for the task taxonomy: (1) Distribution shift to detect: the task focuses on detecting covariate shift or semantic shift; (2) ID data type: the ID data contains one single class or multiple classes; (3) Whether the task requires ID classification; (4) Transductive learning task requires all observations; inductive tasks follow the train-test scheme. Note that OOD detection can encompass a broader spectrum of learning tasks, and go beyond multi-class classification.
Figure 2: Exemplar problem settings for tasks under generalized OOD detection framework. Tags on test images refer to model’s expected predictions. (a) In sensory anomaly detection, test images with covariate shift will be considered as OOD. No semantic shift occurs in this setting. (b) In semantic anomaly detection and one-class novelty detection, normality/ID images belong to one class. Test images with semantic shift will be considered as OOD. No covariate shift occurs in this setting. (c) In multi-class novelty detection, ID images belong to multiple classes. Test images with semantic shift will be considered as OOD. No covariate shift occurs in this setting. (d) Open set recognition is identical to multi-class novelty detection in the task of detection, with the only difference that open set recognition further requires in-distribution classification. (e) Out-of-distribution detection is a super-category that covers semantic AD, one-class ND, multi-class ND, and open-set recognition, which canonically aims to detect test samples with semantic shift without losing the ID classification accuracy. (f) Outlier detection does not follow a train-test scheme. All observations are provided. It fits in the generalized OOD detection framework by defining the majority distribution as ID. Outliers can have any distribution shift from the majority samples.
@@
The detection of semantic distribution shift (e.g., due to the occurrence of new classes) is the focal point of OOD detection tasks considered in this paper, where the label space $\mathcal{Y}$ can be different between ID and OOD data and hence the model should not make any prediction. In addition to OOD detection, several problems adopt the “open-world” assumption and have a similar goal of identifying OOD examples. These include outlier detection (OD) [13, 14, 15, 16], anomaly detection (AD) [17, 18, 19, 20], novelty detection (ND) [21, 22, 23, 24], and open set recognition (OSR) [25, 26, 27]. While all these problems are related to each other by sharing similar motivations, subtle differences exist among the sub-topics in terms of the specific definition. However, existing studies often misuse terms and even datasets, due to a lack of comprehensive understanding of the relations among different problems.
In this survey, we for the first time clarify the similarities and differences between these problems and present a unified framework termed generalized OOD detection. Under this framework, the five problems (i.e., AD, ND, OSR, OOD detection, and OD) can be viewed as special cases or sub-topics. We further conduct a thorough review of each sub-topic and summarize recent technical developments. To sum up, we make three contributions to the OOD detection community:
@@
    A Unified Framework: For the first time, we systematically review five closely related topics of AD, ND, OSR, OOD detection, and OD, and present a more unified framework of generalized OOD detection. Under this framework, the similarities and differences of the five sub-topics can be compared and analyzed. We hope our unification helps the community better understand these problems and correctly position their research in the literature.
@@
    A Comprehensive Survey: We conduct a thorough review of the existing methods developed for each sub-topic, with a particular focus on computer vision and deep learning-based approaches. Despite targeting different problem settings, the methods developed within each area can be generally categorized into four groups: 1) density-based methods, 2) reconstruction-based methods, 3) classification-based methods, and 4) distance-based methods. We hope our survey can help readers build a better understanding of the developments for each problem.
@@
    Future Research Directions: Finally, we draw readers’ attention to some problems or limitations that remain in the current generalized OOD detection field. We conclude this survey with discussions on open challenges and opportunities for future research.
@@ 2 Generalized OOD Detection
Framework Overview  In this section, we introduce a unified framework termed generalized OOD detection, which encapsulates five related sub-topics: anomaly detection (AD), novelty detection (ND), open set recognition (OSR), out-of-distribution detection (OOD), and outlier detection (OD). These sub-topics can be similar in the sense that they all define a certain in-distribution, with the common goal of detecting out-of-distribution samples under the open-world assumption. However, subtle differences exist among the sub-topics in terms of the specific definition and properties of ID and OOD data—which are often overlooked by the research community. To this end, we provide a clear introduction and description of each sub-topic in respective subsections (from Section 2.1 to 2.5). Each subsection details the motivation, background, formal definition, as well as relative position within the unified framework. Applications and benchmarks are also introduced, with concrete examples that facilitate understanding. Thereafter, we conclude this section by discussing and articulating the relationships among the sub-topics (Section 2.6). We also note that AD contains 2 sub-tasks of sensory AD and semantic AD, ND contains 2 sub-tasks of one-class ND and multi-class ND. Therefore, we consider a total of 7 sub-tasks under the framework. Fig. 2 illustrates the settings for each sub-topic.
Preliminary  Key to our framework, the notion of distribution shift is very broad and can exhibit in various forms. There are two general types of distribution shift: covariate shift and semantic (label) shift. Formally, let $\mathcal{X}$ and $\mathcal{Y}$ be the input (sensory) and the label (semantic) space, respectively. A data distribution is defined as a joint distribution $P(X,Y)$ over $\mathcal{X}\times\mathcal{Y}$. Distribution shift can occur in either the marginal distribution $P(X)$, $P(Y)$, or both.
Examples of covariate distribution shift on $P(X)$ include adversarial examples [28, 29], domain shift[30], and style changes[31]. Importantly, we note that covariate shifts are more commonly used to evaluate model generalization and robustness performance, where the label space $\mathcal{Y}$ remains the same during test time. On the other hand, the detection of semantic distribution shift (e.g., due to the occurrence of new classes) is the focal point of many detection tasks considered in this framework, where the label space $\mathcal{Y}$ can be different between ID and OOD data and hence the model should not make any prediction.
With the concept of distribution shift in mind, readers can get a general idea of the differences and connections among sub-topics/tasks in Fig. 1. Next, we proceed with elaborating on each sub-topic.
@@ 2.1 Anomaly Detection
Background  The notion of “anomaly” stands in contrast with the “normal” defined in advance. For example, to create a “hotdog/not-hotdog detector”, we define the concept of hotdog to be normal, and those that violate this definition are identified as anomalies, i.e., not-hotdog. Note that in this case, hotdog is a homogeneous concept, regardless of the possible differences in size, sauce, bread type, price, the origin of the sausage, etc. Any other object, such as steaks, rice, and non-food objects like cats and dogs, are all considered as anomalies.
Note that existing anomaly detection settings often restrict the environment of interest to some specific scenarios. For example, the “hotdog/not-hotdog detector” only focuses on realistic images, assuming the nonexistence of images from domains such as cartoons and sketches. Another realistic example is industrial defect detection, which is based on only one set of assembly lines for a specific product. In other words, the “open-world” assumption is usually not completely “open”. Nevertheless, “not-hotdog” or “defects” can form a large unknown space that breaks the “closed-world” assumption.
In summary, in anomaly detection settings, the key is to treat normality homogeneously, and detect all possible anomalous samples under some specific scenarios.
Definition  Anomaly detection (AD) aims to detect any anomalous samples that are deviated from the predefined normality during testing. The deviation can happen due to either covariate shift or semantic shift, while assuming the other distribution shift do not exist. This leads to two sub-tasks: sensory AD and semantic AD [17], respectively.
Sensory AD detects test samples with covariate shift, under the assumption that normalities come from the same covariate distribution. No semantic shift takes place in sensory AD settings. On the other hand, semantic AD detects test samples with label shift, assuming that normalities come from the same semantic distribution (category), i.e., normalities should belong to only one class. No covariate shift happens in semantic AD settings.
Two broad categories of anomaly detection techniques exist. In the standard unsupervised AD setting, all given training samples are normal samples. The (semi-)supervised AD setting requires a dataset that has been labeled as “normal” and “abnormal”, and involves training a model explicitly.
Position in Framework  Under the generalized OOD detection framework, the definition of “normality” seamlessly connects to the notion of “in-distribution”, and “anomaly” corresponds to “out-of-distribution”. Importantly, AD assumes ID only contains samples from a single class or with a homogeneous characteristic.
@@
Application and Benchmark  Sensory AD only focuses on objects with the same or similar semantics, and identifies the observational differences on their surface. Samples with sensory differences are recognized as sensory anomalies. Example applications include adversarial defense [32], forgery recognition of biometrics and artworks [33, 34, 35, 36], image forensics [37, 38, 39], industrial inspection [40, 41, 42], etc. One popular real-world AD benchmark is MVTec [40] for industrial inspection.
In contrast to sensory AD, semantic AD only focuses on the semantic shift, where covariate shift does not exist. An example of real-world applications is crime surveillance [43, 44]. Active image crawlers for a specific category also need semantic AD methods to ensure the purity of the collected images [45]. An example of the academic benchmarks is to recursively use one class from MNIST as ID during training, and ask the model to distinguish it from the rest of the 9 classes during testing.
Even with different settings, many sensory and semantic AD methods are shown to be mutually inclusive in Section 3.
@@ 2.2 Novelty Detection
Background  The word “novel” generally refers to the unknown, new, and something interesting. Although the goal of novelty detection (ND) is similar to that of AD, there are three differences to note: (1) In terms of motivation, novelty detection usually does not perceive “novel” test samples as erroneous, fraudulent, or malicious as AD does, but cherishes them as learning resources for potential future use with a positive learning attitude [18]; (2) Novelty detection primarily focuses on semantic shift, which is also known as “novel class detection”; (3) Novelty detection removes the restriction that the ID samples should belong to only one class. One or multiple classes can appear during training.
@@
Position in Framework  Under the generalized OOD detection framework, ND deals with the setting where OOD samples have semantic shifts but no covariate shift. Notice that since the in-distribution can contain one or more classes, we distinguish two sub-tasks: one-class ND and multi-class ND. Since one-class ND and semantic AD have the same problem definition except for some nuances in motivation (ref. Section 2.2), their solution space is shared and will be presented in Section 3.
@@ 2.3 Open Set Recognition
Background  Machine learning models trained in the closed-world setting can incorrectly classify test samples from unknown classes as one of the known categories with high confidence [51]. Some literature refers to this notorious overconfident behavior of the model as “arrogance”, or “agnostophobia” [52]. Open set recognition (OSR) is proposed to address this problem, with their own terminology of “known known classes” to represent the categories that exist at training, and “unknown unknown classes” for test categories that do not fall into any training category.
Definition  OSR requires the multi-class classifier to simultaneously: (1) accurately classify test samples that from “known known classes”, and (2) detect test samples from “unknown unknown classes”.
Position in Framework  OSR well aligns with our generalized OOD detection framework, where “known known classes” and “unknown unknown classes” correspond to ID and OOD respectively. Formally, OSR deals with the case where OOD samples during testing have semantic shift, i.e., $P(Y)\neq P^{\prime}(Y)$, but no intentional covariate shift. The goal of OSR is largely shared with that of multi-class ND—the only difference is that OSR additionally requires accurate classification of ID samples from $P(Y)$. We will introduce the methodologies for OSR and multi-class ND together in Section 4.

### v2

@@ Abstract
Out-of-distribution (OOD) detection is critical to ensuring the reliability and safety of machine learning systems. For instance, in autonomous driving, we would like the driving system to issue an alert and hand over the control to humans when it detects unusual scenes or objects that it has never seen during training time and cannot make a safe decision. The term, OOD detection, first emerged in 2017 and since then has received increasing attention from the research community, leading to a plethora of methods developed, ranging from classification-based to density-based to distance-based ones. Meanwhile, several other problems, including anomaly detection (AD), novelty detection (ND), open set recognition (OSR), and outlier detection (OD), are closely related to OOD detection in terms of motiva- tion and methodology. Despite common goals, these topics develop in isolation, and their subtle differences in definition and problem set- ting often confuse readers and practitioners. In this survey, we first present a unified framework called generalized OOD detection, which encompasses the five aforementioned problems, i.e., AD, ND, OSR, OOD detection, and OD. Under our framework, these five problems can be seen as special cases or sub-tasks, and are easier to distinguish. We then review each of these five areas by summarizing their recent technical developments, with a special focus on OOD detection methodologies. We conclude this survey with open challenges and potential research directions.
@@
Out-of-Distribution Detection, Open Set Recognition, Anomaly Detection, Novelty Detection, Outlier Detection
@@ 1 Introduction
Atrustworthy visual recognition system should not only produce accurate predictions on known context, but also detect unknown examples and reject them (or hand them over to human users for safe handling) [1, 2, 3, 4, 5, 6, 7, 8]. For instance, a well-trained food classifier should be able to detect non-food images such as selfies uploaded by users, and reject such input instead of blindly classifying them into existing food categories. In safety-critical applications such as autonomous driving, the driving system must issue a warning and hand over the control to drivers when it detects unusual scenes or objects it has never seen during training.
Figure 1: Taxonomy of generalized OOD detection framework, illustrated by classification tasks. Four bases are used for the task taxonomy: 1) Distribution shift to detect: the task focuses on detecting covariate shift or semantic shift; 2) ID data type: the ID data contains one single class or multiple classes; 3) Whether the task requires ID classification; 4) Transductive learning task requires all observations; inductive tasks follow the train-test scheme. Note that ND is often interchangeable with AD, but ND is more concerned with semantic anomalies. OOD detection is generally interchangeable with OSR for classification tasks.
Figure 2: Illustration of sub-tasks under generalized OOD detection framework with vision tasks. Tags on test images refer to model’s expected predictions. (a) In sensory anomaly detection, test images with covariate shift will be considered as OOD. No semantic shift occurs in this setting. (b) In one-class novelty detection, normal/ID images belong to one class. Test images with semantic shift will be considered as OOD. (c) In multi-class novelty detection, ID images belong to multiple classes. Test images with semantic shift will be considered as OOD. Note that (b) and (c) compose novelty detection, which is identical to the topic of semantic anomaly detection. (d) Open set recognition is identical to multi-class novelty detection in the task of detection, with the only difference that open set recognition further requires ID classification. Out-of-distribution detection solves the same problem as open-set recognition. It canonically aims to detect test samples with semantic shift without losing the ID classification accuracy. However, OOD Detection encompasses a broader spectrum of learning tasks and solution space. (e) Outlier detection does not follow a train-test scheme. All observations are provided. It fits in the generalized OOD detection framework by defining the majority distribution as ID. Outliers can have any distribution shift from the majority.
@@
The detection of semantic distribution shift (e.g., due to the occurrence of new classes) is the focal point of OOD detection tasks, where the label space $\mathcal{Y}$ can be different between ID and OOD data and hence the model should not make any prediction. In addition to OOD detection, several problems adopt the “open-world” assumption and have a similar goal of identifying OOD examples. These include outlier detection (OD) [16, 17, 18, 19], anomaly detection (AD) [20, 21, 22, 23], novelty detection (ND) [24, 25, 26, 27], and open set recognition (OSR) [28, 29, 30]. While all these problems are related to each other by sharing similar motivations, subtle differences exist among the sub-topics in terms of the specific definition. However, the lack of a comprehensive understanding of the relationship between the different sub-topics leads to confusion for both researchers and practitioners. Even worse, these sub-topics, which are supposed to be compared and learned from each other, are developing in isolation.
In this survey, we for the first time clarify the similarities and differences between these problems, and present a unified framework termed generalized OOD detection. Under this framework, the five problems (i.e., AD, ND, OSR, OOD detection, and OD) can be viewed as special cases or sub-topics. We further conduct a literature review for each sub-topic, with a special focus on the OOD detection task. To sum up, we make three contributions to the research community:
@@
    A Unified Framework: For the first time, we systematically review five closely related topics of AD, ND, OSR, OOD detection, and OD, and present a unified framework of generalized OOD detection. Under this framework, the similarities and differences of the five sub-topics can be systematically compared and analyzed. We hope our unification helps the community better understand these problems and correctly position their research in the literature.
@@
    A Comprehensive Survey for OOD Detection: Notice the existence of comprehensive surveys on AD, ND, OSR, and OD methodologies in recent years [20, 21, 22, 23, 29], this survey highlights the overview of OOD detection methods and thus complements existing surveys. For completeness, methodologies of other sub-topics are also reviewed briefly. We hope our survey can help readers build a better understanding of the developments for each problem and their connections.
@@
    Future Research Directions: We draw readers’ attention to some problems or limitations that remain in the current generalized OOD detection field. We conclude this survey with discussions on open challenges and opportunities for future research.
@@ 2 Generalized OOD Detection
Framework Overview  In this section, we introduce a unified framework termed generalized OOD detection, which encapsulates five related sub-topics: anomaly detection (AD), novelty detection (ND), open set recognition (OSR), out-of-distribution (OOD) detection, and outlier detection (OD). These sub-topics can be similar in the sense that they all define a certain in-distribution, with the common goal of detecting out-of-distribution samples under the open-world assumption. However, subtle differences exist among the sub-topics in terms of the specific definition and properties of ID and OOD data—which are often overlooked by the research community. To this end, we provide a clear introduction and description of each sub-topic in respective subsections (from Section 2.1 to 2.5). Each subsection details the motivation, background, formal definition, as well as relative position within the unified framework. Applications and benchmarks are also introduced, with concrete examples that facilitate understanding. Fig. 2 illustrates the settings for each sub-topic. In the end, we conclude this section by introducing the neighborhood topics to clarify the scope of the generalized OOD detection framework. (Section 2.6).
Preliminary: Distribution Shift  Key to our framework, the notion of distribution shift is very broad and can exhibit in various forms. There are two general types of distribution shift: covariate shift and semantic (label) shift. Formally, let $\mathcal{X}$ be the input (sensory) space and $\mathcal{Y}$ be the label (semantic) space, a data distribution is defined as a joint distribution $P(X,Y)$ over $\mathcal{X}\times\mathcal{Y}$. Distribution shift can occur in either the marginal distribution $P(X)$, or both $P(Y)$ and $P(X)$. Note that shift in $P(Y)$ naturally triggers shift in $P(X)$.
Examples of covariate distribution shift on $P(X)$ include adversarial examples [31, 32], domain shift[33], and style changes[34]. Importantly, we note that covariate shift is more commonly used to evaluate model generalization and robustness performance, where the label space $\mathcal{Y}$ remains the same during test time. On the other hand, the detection of semantic distribution shift (e.g., due to the occurrence of new classes) is the focal point of many detection tasks considered in this framework, where the label space $\mathcal{Y}$ can be different between ID and OOD data and hence the model should not make any prediction.
With the concept of distribution shift in mind, readers can get a general idea of the differences and connections among sub-topics/tasks in Fig. 1. Notice that different sub-tasks can be easily identified with the following four dichotomies: 1) covariate / semantic shift dichotomy; 2) single / multiple class dichotomy; 3) ID classification needed / non-needed dichotomy; 4) inductive / transductive dichotomy. Next, we proceed with elaborating on each sub-topic.
@@ 2.1 Anomaly Detection
Background  The notion of “anomaly” stands in contrast with the “normal” defined in advance. The concept of “normal” should be clear and reflect the real task. For example, to create a “not-hotdog detector”, the concept of the normal should be clearly defined as the hotdog class, i.e., a food category, so that objects that violate this definition are identified as anomalies, which include steaks, rice, and non-food objects like cats and dogs. Ideally, “hotdog” would be regarded as a homogeneous concept, regardless of the sub-classes of French or American hotdog.
Current anomaly detection settings often restrict the environment of interest to some specific scenarios. For example, the “not-hotdog detector” only focuses on realistic images, assuming the nonexistence of images from other domains such as sketches. Another realistic example is industrial defect detection, which bases on only one set of assembly lines for a specific product. In other words, the “open-world” assumption is usually not completely “open”. Nevertheless, “not-hotdog” or “defects” can form a large unknown space that breaks the “closed-world” assumption.
In summary, the key to anomaly detection is to define normal clearly (usually without sub-classes) and detect all possible anomalous samples under some specific scenarios.
Definition  Anomaly detection (AD) aims to detect any anomalous samples that are deviated from the predefined normality during testing. The deviation can happen due to either covariate shift or semantic shift, which leads to two sub-tasks: sensory AD and semantic AD, respectively [20].
Sensory AD detects test samples with covariate shift, under the assumption that normalities come from the same covariate distribution. No semantic shift takes place in sensory AD settings. On the other hand, semantic AD detects test samples with label shift, assuming that normalities come from the same semantic distribution (category), i.e., normalities should belong to only one class.
@@
Remark: Sensory / Semantic Dichotomy  Our sensory / semantic dichotomy for the AD sub-task definition comes from the low-level sensory anomalies and high-level semantic anomalies that are introduced in  [35] and highlighted in the recent AD survey [20], to reflect the rise of deep learning. Note that although most sensory and semantic AD methods are shown to be mutually inclusive due to the common shift on $P(X)$, some approaches are specialized in one of the sub-tasks (ref. Section 5). Recent research communities are also trending on subdividing types of anomalies in order to develop targeted methods, so that practitioners can select the optimal solution for their own practical problem [35, 36].
Position in Framework  Under the generalized OOD detection framework, the definition of “normality” seamlessly connects to the notion of “in-distribution”, and “anomaly” corresponds to “out-of-distribution”. Importantly, AD treats ID samples as a whole, which means that regardless of the number of classes (or statistical modalities) in ID data, AD does not require the differentiation in the ID samples. This feature is an important distinction between AD and other sub-topics such as OSR and OOD detection.
Application and Benchmark  Sensory AD only focuses on objects with the same or similar semantics, and identifies the observational differences on their surface. Samples with sensory differences are recognized as sensory anomalies. Example applications include adversarial defense [37], forgery recognition of biometrics and artworks [38, 39, 40, 41], image forensics [42, 43, 44], industrial inspection [45, 46, 47], etc. The most popular academic AD benchmark is MVTec-AD [45] for industrial inspection.
In contrast to sensory AD, semantic AD only focuses on the semantic shift. An example of real-world applications is crime surveillance [48, 49]. Active image crawlers for a specific category also need semantic AD methods to ensure the purity of the collected images [50]. An example of the academic benchmarks is to recursively use one class from MNIST as ID during training, and ask the model to distinguish it from the rest of the 9 classes during testing.
@@
Remark: Alternative Taxonomy on Anomalies  Some previous literature considers anomalies types to be three-fold: point anomalies, conditional or contextural anomalies, and group or collective anomalies [21, 23, 20]. In this survey, we mainly focus on point anomalies detection for its popularity in practical applications and its adequacy to elucidate the similarities and differences between sub-tasks. Details of the other two kinds of anomalies, i.e., contextural anomalies that often occur in time-series tasks, and collective anomalies that are common in the data mining field, are not covered in this survey. We recommend readers to the recent AD survey papers [20] for in-depth discussion on them.
Remark: Taxonomy based on Supervision  We use sensory / semantic dichotomy is to subdivide AD at the task level. From the perspective of methodologies, some literature categorizes AD techniques into unsupervised and (semi-) supervised settings. Note that these two taxonomies are orthogonal as they focus on tasks and methods respectively.
@@ 2.2 Novelty Detection
Background  The word “novel” generally refers to the unknown, new, and something interesting. While novelty detection (ND) is often interchangeable with AD in the community, strictly speaking, their subtle difference is worth noticing. In terms of motivation, novelty detection usually does not perceive “novel” test samples as erroneous, fraudulent, or malicious as AD does, but cherishes them as learning resources for potential future use with a positive learning attitude [21, 20]. In fact, novelty detection is also known as “novel class detection” [26, 27], indicating that it is primarily focusing on detecting semantic shift.
@@
Position in Framework  Under the generalized OOD detection framework, ND deals with the setting where OOD samples have semantic shift, without the need for classification in the ID set even if possible. Therefore, ND shares the same problem definition with semantic AD.
@@
Remark: One-Class / Multi-Class Dichotomy  Although the ND models do not require the ID classification even with multi-class annotations, the method on multi-class ND can be different from one-class ND, as multi-class ND can make use of the multi-class classifier while one-class ND cannot. Also note that semantic AD can be further split into one-class semantic AD and multi-class semantic AD that matches ND, as semantic AD is equivalent to ND.
Remark: Nuance between AD and ND  Apart from the special interest in semantics, some literature [56, 57] also point out that ND is supposed to be fully unsupervised, while AD might have some abnormal training samples.
@@ 2.3 Open Set Recognition
Background  Machine learning models trained in the closed-world setting can incorrectly classify test samples from unknown classes as one of the known categories with high confidence [58]. Some literature refers to this notorious overconfident behavior of the model as “arrogance”, or “agnostophobia” [59]. Open set recognition (OSR) is proposed to address this problem, with their own terminology of “known known classes” to represent the categories that exist at training, and “unknown unknown classes” for test categories that do not fall into any training category. Some other terms, such as open category detection [60] and open set learning [61], are simply different expressions for OSR.
Definition  Open set recognition requires the multi-class classifier to simultaneously: 1) accurately classify test samples from “known known classes”, and 2) detect test samples from “unknown unknown classes”.
Position in Framework  OSR well aligns with our generalized OOD detection framework, where “known known classes” and “unknown unknown classes” correspond to ID and OOD respectively. Formally, OSR deals with the case where OOD samples during testing have semantic shift, i.e., $P(Y)\neq P^{\prime}(Y)$. The goal of OSR is largely shared with that of multi-class ND—the only difference is that OSR additionally requires accurate classification of ID samples from $P(Y)$.